A ball is thrown upward with an initial velocity of 64 ft/sec from a height 960 ft. Its height h, in feet, after t seconds is given by h(t)=-16t2 +64t+960. After how long will the ball reach the ground ?
Answer by lwsshak3(11628)

A ball is thrown upward with an initial velocity of 64 ft/sec from a height 960 ft. Its height h, in feet, after t seconds is given by h(t)=-16t2 +64t+960. After how long will the ball reach the ground ?
***
ball will reach the ground when (h(t)=0
-16t2 +64t+960=0
-t^2+4t+60=0
(-t+10)(t+6)=0
t=-6(reject)
or
t=10
After how long will the ball reach the ground ? 10 sec